OpenStudy (anonymous):

Lim x-->0 sqrt(4+x) -2/ x

OpenStudy (anonymous):

\[\lim_{x \rightarrow o}\frac{ \sqrt{4+x} -2 }{ x }=?\]

OpenStudy (anonymous):

try conjugate

OpenStudy (anonymous):

you can differenciate it and then put the value in the result

OpenStudy (anonymous):

rationalise the numerator.

OpenStudy (anonymous):

i think the asnwer is 1/4

OpenStudy (anonymous):

I have the answer but don't understand how to work it out. The answer is 1/4

OpenStudy (anonymous):

when u differrnciate it then

OpenStudy (anonymous):